The diagonal of a square frame is 6 inches. What is the length of each side of the frame?
Answer by mananth(16949)

the diagonal is the hypotenuse.
the legs are the sides of the square.
6^2=x^2+x^2
36=2x^2
18=x^2
x=sqrt(18)
x= 4.24 in